An electric induction coil of length of 10cm has 1000 turns of radius 5cm. (μ0=4π x 10-7S.I). What is the value of the inductance?


1
Expert's answer
2021-07-29T00:34:02-0400

"L=\frac{\mu_0 N^2S}l=\frac{\mu_0\pi N^2r^2}l=\frac{4\pi^2\cdot 10^{-7}\cdot 1000^2\cdot 0.05^2}{0.1}=0.1~H."